In triangle ABC if r=1 R=3 s= 5 then area of triangle ABC

Grade:11

2 Answers

Arun
6 years ago
 
Dear Pranit
 
Area of triangle (Delta) = s *r 
Area = 5*1 = 5 unit²
 
Regards
Arun (askiitians forum expert)
Rajesh Kurlepu
2 years ago
The question is incorrect since we don't get real values of a,b and c
We know ∆=r×s From this equation we get ∆=5
We also know that 4∆R=abc
From the above equation we get abc=60
On Applying AM-GM inequality for a,b,c we can clearly observe that the inequality doesn't hold good in this case
So no such triangle exists!!
Even if u try solving u will get ab+bc+ca=38
If u solve cubic equation with roots a,b,c then all its roots are not real!!
 
Therefore the question is incorrect